OpenStudy (anonymous):

Find the probability of the following five-card poker hands from a 52-card deck. In poker, aces are either high or low. Two pair (2 cards of one value, 2 of another value)

OpenStudy (anonymous):

Prob of first card is 52/52 --- Prob of second card (pairs first) = 3/51 --- Prob of third card (not same as first pair) = 48/50 --- Prob of fourth (pairs third) = 3/49 --- Prob of fifth (any card not the same as first two pairs)=44/48 multiply these probabilites together= 0.003169

OpenStudy (anonymous):

thank you